# Koopman for stochastic dynamics: error bounds for kernel extended dynamic mode decomposition

### v1

@@ 3. $H^{s}$-invariance under the Koopman operator
Since Wendland (and Matérn) kernels generate the Sobolev spaces as their respective native spaces, we study the invariance of Sobolev spaces $H^{s}$ under the stochastic Koopman operator. As those spaces are defined via partial integration, a condition for preserving weak derivatives under Koopman operator action is paramount. The central ingredient for this is the following signed measure associated with the transition kernel $\rho_{x}$. For $g\in L^{\infty}(\Omega)$, define $\nu_{g}:\mathcal{B}_{\Omega}\to\mathbb{R}$ by
$\nu_{g}(\Delta):=\int_{\Omega}\rho_{x}(\Delta)g(x)\,dx.$ (22)
@@
Assume that ($L^{1}$) holds and consider $\mathcal{K}_{1}:L^{1}(\Omega)\to L^{1}(\Omega)$ as well as $\mathcal{K}_{2}:L^{2}(\Omega)\to L^{2}(\Omega)$. Then, the measure $\nu_{g}$ defined in (22) is absolutely continuous w.r.t. the Lebesgue measure, and its density $\frac{d\nu_{g}}{dx}\in L^{\infty}(\Omega)$ satisfies
$\mathcal{K}_{1}^{*}g=\mathcal{K}_{2}^{*}g=\frac{d\nu_{g}}{dx},\quad\mathrm{for\ all}\quad g\in L^{\infty}(\Omega).$
@@ 3.1. Conditions for $H^{s}$-invariance and boundedness
Assume that the condition ($L^{1}$) holds.
@@
    Let $\ell\in\mathbb{N}_{\geq 1}$. Then $H^{\ell}$ is invariant under $\mathcal{K}$ if and only if there is $C\geq 0$ such that
@@
    If (23) holds, then $H^{s}$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
@@
    If for some constant $C>0$ and $\ell\in\mathbb{N}_{\geq 1}$
    then $H^{s}$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
@@
For $n=1$ and $\Omega=(0,1)\subset\mathbb{R}$, assume $Y\mid\{X=x\}\sim\rho_{x}:=\mathcal{N}_{(0,1)}(x,1)$ where $\mathcal{N}_{(0,1)}(x,1)$ denotes the (doubly) truncated normal distribution on $(0,1)$ with location parameter $x\in(0,1)$ and scale parameter set to $1$. Then, $H^{1}(\Omega)$ is invariant under the associated Koopman operator that acts on $f\in H^{1}(\Omega)$ via $f\in H^{1}\bigl((0,1)\bigr)$ via $\mathcal{K}f(x)=\mathbb{E}[f(Y)\mid X=x]$.
@@
The doubly truncated normal distribution $\rho_{x}$ is the distribution of a random variable $Z\sim\mathcal{N}(x,1)$ conditioned on the event $0<Z<1$. To apply part (iii) of Theorem 3.2, we study. the density of $\rho_{x}$. Denote the cumulative distribution function of the standard normal distribution by $F(\xi)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\xi}e^{-t^{2}/2}\,dt$. Then, the distribution of $Y$ conditioned on $\{X=x\}$ with $x\in(0,1)$ has density $p_{x}=p(x,\cdot):(0,1)\to[0,\infty)$ given by
@@
Numerator and denominator both depend smoothly on $x$, so we can easily compute derivatives with respect to $x$ and bound them. Using $F^{\prime}(\xi)=\frac{1}{\sqrt{2\pi}}e^{-\xi^{2}/2}$, we get
$\partial_{1}p(x,y)=\frac{\partial_{1}G(x,y)H(x)-G(x,y)H^{\prime}(x)}{H(x)^{2}}=G(x,y)\left((y-x)-\frac{H^{\prime}(x)}{H(x)^{2}}\right)$
Note that $0<C_{1}\leq H(x)\leq C_{2}$ since $H(x)=\mathbb{P}(U\in(-x,1-x))$ for some $U\sim\mathcal{N}(0,1)$ and $x\in(0,1)$. Using the fact that $G(x,y)\leq\frac{1}{\sqrt{2\pi}}$ for all $x,y\in(0,1)$ and $|H^{\prime}(x)|\leq\frac{1}{\sqrt{2\pi}}$, we obtain the basic bound
@@ 3.2. Sobolev Spaces
For self-containedness, we briefly recapitulate the notion of Sobolev spaces following [Adams2003]. Readers who are familiar with Sobolev spaces may only skim this section for notation or skip it entirely.
Recall the bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary and write $C_{c}^{\infty}(\Omega)$ for the space of infinitely differentiable and compactly supported functions on $\Omega$. If the domain is clear from context, we might write $C_{c}^{\infty}=C_{c}^{\infty}(\Omega)$. Given some multi-index $\alpha\in\mathbb{N}_{0}^{n}$, we write $D^{\alpha}\varphi$ for the $\alpha$-th partial derivative of a sufficiently smooth function $\varphi$. Note that $\varphi\in C_{c}^{\infty}(\Omega)$ implies $D^{\alpha}\varphi(x)=0$ for all $x\in\partial\Omega$ and $\alpha\in\mathbb{N}^{n}$ since $\varphi\equiv 0$ in an open neighborhood of the boundary.
Next, recall the notion of the weak derivative of a function: Consider a locally integrable function $f\in L_{\mathrm{loc}}^{1}(\Omega)$; that is, $f$ is integrable on every compact set $K\subset\Omega$, for which $f\in L^{2}(\Omega)$ is clearly sufficient. If there exists a function $g\in L_{\mathrm{loc}}^{1}(\Omega)$ satisfying
@@
for every test function $\varphi\in C_{c}^{\infty}(\Omega)$, we call $g$ a $\alpha$-th weak derivative of $f$. The weak derivative is unique up to a set of measure zero, i.e., two weak derivatives of $f$ agree almost everywhere, and needs not be defined pointwise (the elements of $L_{\mathrm{loc}}^{1}(\Omega)$ are equivalence classes, after all). In the case of a “conventionally” partially differentiable function, the weak derivative coincides (almost everywhere) with the respective partial derivative. Hence, in a slight abuse of notation, we write $g=D^{\alpha}f$ for the weak derivative.
Using the notion of a weak derivative, given some $\ell\in\mathbb{N}$ we introduce
@@
Sobolev spaces can be generalized to real-valued regularity order $s\geq 0$. Those fractional Sobolev spaces arise, for example, as reproducing kernel Hilbert spaces of certain kernels. There are two predominant definitions of fractional Sobolev spaces in the literature, either via the Fourier transform or via weak derivatives and interpolation between integer-order Sobolev spaces. For a domain with Lipschitz boundary, as is $\Omega\subset\mathbb{R}^{n}$, those definitions coincide [McLe2000, Theorem 3.30], so we proceed with the more suitable second definition for the remainder of this paper. Given $s\geq 0$, write $s=\lfloor s\rfloor+r$ with $r\in[0,1)$ and define the norm
@@
Under this definition, the Sobolev space $H^{s}(\Omega)=\{f\in L^{2}(\Omega):\|f\|_{H^{s}(\Omega)}<\infty\}$ presents itself as the interpolation space between $H^{\lfloor s\rfloor}(\Omega)$ and $H^{\lceil s\rceil}(\Omega)$.
@@
We start by proving the important auxiliary result from the beginning of this section. While Lemma 3.1 essentially follows from [PhilScha24, Appendix B], we provide a concise proof below for self-containedness.
@@
Let $g\in L^{\infty}(\Omega)$ and $\Delta\in\mathcal{B}_{\Omega}$. If ($L^{1}$) holds, the definition of (22) immediately gives
$|\nu_{g}(\Delta)|\,\leq\,\|g\|_{L^{\infty}}\int\rho_{x}(\Delta)\,dx\,\leq\,C\|g\|_{L^{\infty}}|\Delta|.$
@@
$\langle\mathcal{K}_{1}f,g\rangle_{L^{1},L^{\infty}}=\int\mathcal{K}_{1}f\cdot g\,dx=\int\int f\,d\rho_{x}g(x)\,dx=\int f\,d\nu_{g}=\left\langle f,\frac{d\nu_{g}}{dx}\right\rangle_{L^{1},L^{\infty}}.$
This shows $\mathcal{K}_{1}^{*}g=\frac{d\nu_{g}}{dx}$. Finally, if $f,g\in L^{\infty}(\Omega)$, we have
@@
$\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-1}}=\|\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\|_{H^{-1}}$ $=\sup\left\{|\langle f,\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\rangle_{H^{1},H^{-1}}|:f\in H^{1},\,\|f\|_{H^{1}}=1\right\}\,$
$\leq\,C_{1}\|\varphi\|_{L^{2}}.$
@@
for every $|\alpha|=1$. Hence, $|\langle\mathcal{K}f,D^{\alpha}\varphi\rangle|=|\langle f,K_{2}^{\ast}(D^{\alpha}\varphi)\rangle|\leq C_{\alpha}\|f\|_{H^{1}}\|\varphi\|_{L^{2}}$ holds for all $f\in H^{1}$ and all $\varphi\in C_{c}^{\infty}$. Therefore, for fixed $f\in H^{1}$, the functional $\varphi\mapsto\langle\mathcal{K}f,D^{\alpha}\varphi\rangle$ is bounded on $C_{c}^{\infty}$ and can thus be extended to a bounded functional $F_{f}$ on $L^{2}$. By Riesz’ representation theorem, $F_{f}\varphi=\langle Lf,\varphi\rangle$ for all $\varphi\in L^{2}$ with some $Lf\in L^{2}$. Equating those expressions gives $\langle\mathcal{K}f,D^{\alpha}\varphi\rangle=F_{f}\varphi=\langle Lf,\varphi\rangle$ for all $f\in H^{1}$ and all $\varphi\in C_{c}^{\infty}$. This shows that $\mathcal{K}f$ has the weak derivative $Lf=D^{\alpha}(\mathcal{K}f)\in L^{2}$ for all $f\in H^{1}$ and all $|\alpha|=1$, hence $\mathcal{K}f\in H^{1}$ and $H^{1}$ is invariant under $\mathcal{K}$. This shows the equivalence from (i) in Theorem 3.2.
Step 2: $H^{\ell}$-invariance, $\ell\in\mathbb{N}$. The characterization from step 1 is easily generalized to Sobolev spaces of higher smoothness. Fix $\ell\in\mathbb{N}$. By definition, we have $f\in H^{\ell}$ if and only if $D^{\alpha}f\in L^{2}$ for every $\alpha\in\mathbb{N}^{n}$ with $|\alpha|\leq\ell$. For the sufficiency, assume the invariance of $H^{\ell}$, i.e., $\mathcal{K}f\in H^{\ell}$ for every $f\in H^{\ell}$. Again, this gives the boundedness of $\mathcal{K}:H^{\ell}\to H^{\ell}$ by Proposition 2.4, and for all $\alpha\in\mathbb{N}^{n}$ with $|\alpha|\leq\ell$ we have $D^{\alpha}(\mathcal{K}f)\in L^{2}$. By the definition of the weak derivative (25)
@@
Step 3: $H^{s}$-invariance, $s\in(0,\ell]$. By [BrenScot2008, Theorem 14.2.3], for real-valued $s\in(0,\ell]$, the space $H^{s}$ coincides with the real interpolation space
$H^{s}(\Omega)=[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta},\quad\theta=s/\ell\in(0,1)$
with equivalent norms. Under condition ($L^{1}$), we have the invariance of $L^{2}(\Omega)$ under $\mathcal{K}$. Let (23) hold for $\ell\in\mathbb{N}$. Then $H^{\ell}(\Omega)$ is also invariant under $\mathcal{K}$ by step 2. [Tart2007, Theorem 22.3] now implies that the Koopman operator
@@ 4. Error bounds for approximations of Koopman operators of stochastic systems
To verify especially the condition $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ in the assumption above, we refer to the sufficient conditions of the previous Section 3, where Theorem 3.2 guarantees the invariance of $\mathbb{H}$ since Sobolev spaces $\mathbb{H}=H^{s}$ are induced by Wendland kernels as RKHS [KohnPhil24, Theorem 4.1].
We start by constructing the kernel-based approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ of the stochastic Koopman operator from (18) more thoroughly. In an intermediate step, we derive a kernel EDMD approximant $\widehat{\mathcal{K}}$, significantly building upon the analogue from [KohnPhil24] in the deterministic setting. This allows to split the approximation error as follows:
@@
The remainder of this section works towards the proof of Theorem 2.6 by analyzing both errors independently. The intermediary approximation $\widehat{\mathcal{K}}$ is employed in Section 4.1 to prove a bound on the deterministic projection error, see Theorem 4.2. Therein, we appeal to similar arguments as [KohnPhil24], where the invariance of a reproducing kernel Hilbert space under the Koopman operator is assumed as in our Assumption 4.1. Since $\widehat{\mathcal{K}}$ requires the exact knowledge of propagated observables under the stochastic Koopman operator, i.e., conditional expectation, Section 4.2 proposes the random, but computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$. The approximation of the “exact”, but not computable $\widehat{\mathcal{K}}$ introduces a probabilistic error that is dealt with in Theorem 4.6.

### v2

@@ 3. $H^{s}$-invariance under the Koopman operator
Since Wendland (and Matérn) kernels generate the Sobolev spaces $H^{s}(\Omega)$ as their respective native spaces, we study their invariance under the stochastic Koopman operator. As those spaces are defined via partial integration, a condition for preserving weak derivatives under Koopman operator action is paramount. The central ingredient for this is the following signed measure associated with the transition kernel $\rho_{x}$. For $g\in L^{\infty}(\Omega)$, define $\nu_{g}:\mathcal{B}_{\Omega}\to\mathbb{R}$ by
$\nu_{g}(\Delta):=\int_{\Omega}\rho_{x}(\Delta)g(x)\,dx,\qquad\Delta\in\mathcal{B}_{\Omega}.$ (23)
@@
Assume that (4) holds and consider $\mathcal{K}_{1}:L^{1}(\Omega)\to L^{1}(\Omega)$ as well as $\mathcal{K}_{2}:L^{2}(\Omega)\to L^{2}(\Omega)$. Then, the measure $\nu_{g}$ defined in (23) is absolutely continuous w.r.t. the Lebesgue measure, and its density $\frac{d\nu_{g}}{dx}\in L^{\infty}(\Omega)$ satisfies
$\mathcal{K}_{1}^{*}g=\mathcal{K}_{2}^{*}g=\frac{d\nu_{g}}{dx}\quad\mathrm{for\ all}\quad g\in L^{\infty}(\Omega),$
@@ 3.1. Conditions for $H^{s}$-invariance and boundedness
Assume that (4) holds.
@@
    Let $\ell\in\mathbb{N}_{\geq 1}$. Then $H^{\ell}(\Omega)$ is invariant under $\mathcal{K}$ if and only if there is $C\geq 0$ such that
@@
    If (24) holds, then $H^{s}(\Omega)$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
@@
    If for some constant $C>0$ and $\ell\in\mathbb{N}_{\geq 1}$, we have
    $\sum_{|\alpha|\leq\ell}\left\|D_{1}^{\alpha}\,p(\cdot,\cdot)\right\|_{L^{2}(\Omega\times\Omega)}\leq C,$ (25)
    then $H^{s}(\Omega)$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
@@
For $n=1$ and $\Omega=(0,1)\subset\mathbb{R}$, assume $Y\mid\{X=x\}\sim\rho_{x}:=\mathcal{N}_{(0,1)}(x,1)$, where $\mathcal{N}_{(0,1)}(x,1)$ denotes the (doubly) truncated normal distribution on $(0,1)$ with location parameter $x\in(0,1)$ and scale parameter set to $1$. Then, $H^{1}(\Omega)$ is invariant under the associated Koopman operator that acts on $f\in H^{1}(\Omega)$ via $f\in H^{1}\bigl((0,1)\bigr)$ via $\mathcal{K}f(x)=\mathbb{E}[f(Y)\mid X=x]$.
@@
The doubly truncated normal distribution $\rho_{x}=\mathcal{N}_{(0,1)}(x,1)$ is defined as the distribution of a random variable $Z\sim\mathcal{N}(x,1)$ conditioned on the event $0<Z<1$. To apply part (iii) of Theorem 3.2, we study the density of the transition kernel $\rho_{x}$. Denote the cumulative distribution function of the standard normal distribution by $F(\xi)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\xi}e^{-t^{2}/2}\,dt$. Then, the distribution of $Y$ conditioned on $\{X=x\}$ with $x\in(0,1)$ has density $p_{x}=p(x,\cdot):(0,1)\to[0,\infty)$ given by
@@
The numerator and denominator both depend smoothly on $x$. With $F^{\prime}(\xi)=\frac{1}{\sqrt{2\pi}}e^{-\xi^{2}/2}$, we compute the derivative with respect to $x$:
$\partial_{1}p(x,y)=\frac{\partial_{1}G(x,y)H(x)-G(x,y)H^{\prime}(x)}{H(x)^{2}}=G(x,y)\left((y-x)-\frac{H^{\prime}(x)}{H(x)^{2}}\right).$
Note that $0<C_{1}\leq H(x)\leq C_{2}$ since $H(x)=\mathbb{P}(U\in(-x,1-x))$ for some $U\sim\mathcal{N}(0,1)$ and $x\in(0,1)$. With $G(x,y)\leq\frac{1}{\sqrt{2\pi}}$ for all $x,y\in(0,1)$ and $|H^{\prime}(x)|\leq\frac{1}{\sqrt{2\pi}}$, we obtain the bound
@@ 3.2. Sobolev Spaces
For self-containedness, we briefly recapitulate the notion of Sobolev spaces following [1]. Readers who are familiar with Sobolev spaces may only skim this section for notation or skip it entirely.
Consider the bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary and the space $C_{c}^{\infty}(\Omega)$ of infinitely differentiable and compactly supported functions on $\Omega$. We recall the notion of the weak derivative of a function: Consider a locally integrable function $f\in L_{\mathrm{loc}}^{1}(\Omega)$; that is, $f$ is integrable on every compact set $K\subset\Omega$, for which $f\in L^{2}(\Omega)$ is clearly sufficient. If there exists a function $g\in L_{\mathrm{loc}}^{1}(\Omega)$ satisfying
@@
for every test function $\varphi\in C_{c}^{\infty}(\Omega)$, we call $g$ the $\alpha$-th weak derivative of $f$ and write $g=D^{\alpha}f$. Using the notion of a weak derivative, given some $\ell\in\mathbb{N}$ we introduce
@@
Sobolev spaces can be generalized to real-valued regularity order $s\geq 0$. Those fractional Sobolev spaces arise, for example, as reproducing kernel Hilbert spaces of certain kernels. There are two predominant definitions of fractional Sobolev spaces in the literature, either via the Fourier transform or via weak derivatives and interpolation between integer-order Sobolev spaces. For a domain with Lipschitz boundary, as is $\Omega\subset\mathbb{R}^{n}$, those definitions coincide by [31, Theorem 3.30], so we proceed with the second definition which is more suitable for the remainder of this paper. Given $s\geq 0$, write $s=\lfloor s\rfloor+r$ with $r\in[0,1)$ and define the norm
@@
Under this definition, the Sobolev space $H^{s}(\Omega)=\{f\in L^{2}(\Omega):\|f\|_{H^{s}(\Omega)}<\infty\}$ is an interpolation space between $H^{\lfloor s\rfloor}(\Omega)$ and $H^{\lceil s\rceil}(\Omega)$. More precisely, $H^{s}(\Omega)=[H^{\lfloor s\rfloor}(\Omega),H^{\lceil s\rceil}(\Omega)]_{r}$.
@@
We start by proving the important auxiliary result from the beginning of this section. While Lemma 3.1 essentially follows from [39, Appendix B], we provide a concise proof below for self-containedness.
@@
Let $g\in L^{\infty}(\Omega)$ and $\Delta\in\mathcal{B}_{\Omega}$. If (4) holds, the definition of (23) immediately gives
$|\nu_{g}(\Delta)|\,\leq\,\|g\|_{L^{\infty}}\int_{\Omega}\rho_{x}(\Delta)\,dx\,\leq\,C\|g\|_{L^{\infty}}|\Delta|.$
@@
$\langle\mathcal{K}_{1}f,g\rangle_{L^{1},L^{\infty}}=\int_{\Omega}\mathcal{K}_{1}f\cdot g\,dx=\int_{\Omega}\int_{\Omega}f\,d\rho_{x}g(x)\,dx=\int_{\Omega}f\,d\nu_{g}=\left\langle f,\frac{d\nu_{g}}{dx}\right\rangle_{L^{1},L^{\infty}}.$
This shows $\mathcal{K}_{1}^{*}g=\frac{d\nu_{g}}{dx}$. Finally, since $|\Omega|<\infty$ implies the inclusion $L^{\infty}(\Omega)\subset L^{2}(\Omega)$, we have for $f,g\in L^{\infty}(\Omega)$:
@@
$\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-1}}=\|\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\|_{H^{-1}}$ $=\sup\left\{\left|\langle f,\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\rangle_{H^{1},H^{-1}}\right|:f\in H^{1},\,\|f\|_{H^{1}}=1\right\}\,$
$\leq\,C_{1}\|\varphi\|_{L^{2}}$
@@
for every $|\alpha|=1$. Hence, $|\langle\mathcal{K}f,D^{\alpha}\varphi\rangle|=|\langle f,K_{2}^{\ast}(D^{\alpha}\varphi)\rangle|\leq C_{\alpha}\|f\|_{H^{1}}\|\varphi\|_{L^{2}}$ holds for all $f\in H^{1}$ and all $\varphi\in C_{c}^{\infty}$. Therefore, for fixed $f\in H^{1}$, the functional $\varphi\mapsto\langle\mathcal{K}f,D^{\alpha}\varphi\rangle$ is bounded on $C_{c}^{\infty}$ and can be extended to a bounded functional $F_{f}$ on $L^{2}$. By Riesz’ representation theorem, $F_{f}\varphi=\langle Lf,\varphi\rangle$ for all $\varphi\in L^{2}$ with some $Lf\in L^{2}$. Equating those expressions gives $\langle\mathcal{K}f,D^{\alpha}\varphi\rangle=F_{f}\varphi=\langle Lf,\varphi\rangle$ for all $f\in H^{1}$ and all $\varphi\in C_{c}^{\infty}$. This shows that $\mathcal{K}f$ has the weak derivative $Lf=D^{\alpha}(\mathcal{K}f)\in L^{2}$ for all $f\in H^{1}$ and all $|\alpha|=1$, hence $\mathcal{K}f\in H^{1}$ and $H^{1}$ is invariant under $\mathcal{K}$. This shows the equivalence from (i) in Theorem 3.2.
Step 2: $H^{\ell}$-invariance, $\ell\in\mathbb{N}$. The characterization from step 1 is easily generalized to Sobolev spaces of higher smoothness. Fix $\ell\in\mathbb{N}$. By definition, we have $f\in H^{\ell}$ if and only if $D^{\alpha}f\in L^{2}$ for every $\alpha\in\mathbb{N}^{n}$ with $|\alpha|\leq\ell$. For the sufficiency, assume the invariance of $H^{\ell}$, i.e., $\mathcal{K}f\in H^{\ell}$ for every $f\in H^{\ell}$. Again, this gives the boundedness of $\mathcal{K}:H^{\ell}\to H^{\ell}$ by Proposition 2.4, and for all $\alpha\in\mathbb{N}^{n}$ with $|\alpha|\leq\ell$ we have $D^{\alpha}(\mathcal{K}f)\in L^{2}$. By the definition of the weak derivative (26), we have
@@
Step 3: $H^{s}$-invariance, $s\in(0,\ell]$. By [8, Theorem 14.2.3], for real-valued $s\in(0,\ell]$, the space $H^{s}$ coincides with the real interpolation space
$H^{s}(\Omega)=[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta},\quad\theta=s/\ell\in(0,1),$
with equivalent norms. Under condition (4), we have the invariance of $L^{2}(\Omega)$ under $\mathcal{K}$. Let (24) hold for $\ell\in\mathbb{N}$. Then $H^{\ell}(\Omega)$ is also invariant under $\mathcal{K}$ by step 2. [49, Theorem 22.3] now implies that the Koopman operator
@@ 4. Error bounds for approximations of Koopman operators of stochastic systems
To verify the condition $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ in Assumption 4.1, we refer to the sufficient conditions of the previous Section 3, where Theorem 3.2 guarantees the invariance of $\mathbb{H}$ since Sobolev spaces $\mathbb{H}=H^{s}$ are induced by Wendland kernels as RKHS [24, Theorem 4.1].
We start by constructing the kernel-based approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ of the stochastic Koopman operator from (19) more thoroughly. In an intermediate step, we derive a kernel EDMD approximant $\widehat{\mathcal{K}}$, significantly building upon the analogue from [24] in the deterministic setting. This allows to split the approximation error as follows:
@@
The remainder of this section works towards the proof of Theorem 2.6 by analyzing both errors independently. The intermediary approximation $\widehat{\mathcal{K}}$ is employed in Section 4.1 to prove a bound on the deterministic projection error, see Theorem 4.2. Therein, we appeal to similar arguments as [24], where the invariance of a reproducing kernel Hilbert space under the Koopman operator is assumed as in our Assumption 4.1. Since $\widehat{\mathcal{K}}$ requires the exact knowledge of propagated observables under the stochastic Koopman operator, i.e., conditional expectations, Section 4.2 proposes the random, but computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$. The approximation of the “exact”, but not computable $\widehat{\mathcal{K}}$ introduces a probabilistic error that is dealt with in Theorem 4.6.